A block of wood floats in water with 0.546 of its volume submerged. In oil it has
0.918 of its volume submerged. Find the density of (a).the wood (b).the oil.

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Dear student 
Let density of wood be x 
For floating in water 
ρbvg = 0.546 * dw * g 
For floating in oil 
ρbvg = 0.918*doil*g 
so, doil/dw = 0.546/0.918 
So, doil / dw = 91/153 
Hope this will help 
Good Luck 
Cheers